OpenStudy (anonymous):

The surface area (A) of a box is A = 2LW + 2LH + 2WH. Find W, if A = 94, L = 3, and H = 5. W = 106 W = 59 W = 4 W = 8

OpenStudy (anonymous):

6w+30+10w = 94 16w = 94-30 16w = 64 w = 64/16 w = 4

OpenStudy (anonymous):

thnx!!!